Here's the question:
Solve the formula for converting celsius to Fahrenheit for C. Solve the formula for converting fahrenheit to Celsius for F. What do you notice?

Answer by TimothyLamb(4379)

linear equation for converting C (x-axis) to F (y-axis):
(0,32)
(100,212)
---
user the slope-intercept form:
y = mx + b
---
m = slope = dy/dx
m = (212-32)/(100-0)
m = 180/100
m = 9/5
---
b = y-intercept (y-coordinate when x=0)
b = 32
F = (9/5)C + 32
---
linear equation for converting F (x-axis) to C (y-axis):
(32,0)
(212,100)
---
user the point-slope form:
y - y1 = m(x + x1)
---
m = slope = dy/dx
m = (100-0)/(212-32)
m = 100/180
m = 5/9
---
C - 0 = 5/9(F - 32)
C = 5/9(F - 32)
---
answer:
F = (9/5)C + 32
C = (5/9)(F - 32)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php